Determine the mass of Na2SO3 necessary to prepare 250.0 mL of a solution that is 0.463 M Na2SO3
Answer #1

volume , V = 2.5*10^2 mL

= 0.25 L

use:

number of mol,

n = Molarity * Volume

= 0.463*0.25

= 0.1158 mol

Molar mass of Na2SO3,

MM = 2*MM(Na) + 1*MM(S) + 3*MM(O)

= 2*22.99 + 1*32.07 + 3*16.0

= 126.05 g/mol

use:

mass of Na2SO3,

m = number of mol * molar mass

= 0.1158 mol * 1.26*10^2 g/mol

= 14.59 g

Answer: 14.6 g